# Thermal condensation with polybaric pressure

### v1

@@ Abstract
In this work, we have carried out a two-dimensional (2D) simulation of thermal instability (TI) in interstellar matter (ISM), considering it to be a weakly ionized inviscid plasma with radiation loss. We carry out the simulation with the help of our multi-fluid flux-corrected transport (mFCT) code which incorporates a background magnetic field and anisotropic pressure. The anisotropic pressure is modeled with a polybaric pressure model (Stasiewicz, 2004). We find that the resultant turbulent state of the interstellar matter (ISM), driven purely by time-dependent heating and cooling with polybaric pressure is similar in nature to the one driven by external forcing such as by supernova feedback. This demonstrates the universal nature of the turbulence in ISM, which is largely isobaric and incompressible. The findings of our analysis is consistent with the contemporary status of knowledge about the multiphase nature of the ISM with volume and mass fractions of the various components of the ISM that is warm, cold, and unstable neutral matter (UNM) in the ranges reported by various numerical and observational analysis.Though the strength of the background magnetic field only marginally affects the overall evolution, the ratio of the parallel and perpendicular pressures can considerably alter the mass and volume fractions of the three phases, which can affect the overall evolution of the TI in the long run.
Magnetohydrodynamics(1964) — Interstellar medium(847) — Interstellar phases(850) —Interstellar dynamics(839)
@@ 1 Introduction
Thermal instability (TI) is a physical process responsible for non-gravitational condensations in various plasma environments such as solar corona, solar wind, magnetosheath, interstellar medium (ISM), intracluster medium (ICM) etc. When thermal equilibrium becomes unstable, it can change the structure and dynamics of the medium at a small scale resulting in condensations of high density and low temperature such as clouds and filaments, known as the precursors to star formation. First introduced by Parker (1953) and then by Field in his seminal work (1965), TI is known to explain the existence of multiphase structure of the ISM and mass flow between different intermediate structures, which has very significant effects on the rate of star formation (Chieze, 1987).
@@
The multiphase model of the ISM is a well-studied physical model, which was thought to be consisting of two stable phases – warm and cold, formed out of neutral atomic hydrogen (Field et al., 1969). Studies have later confirmed the existence of these two phases – diffuse warm neutral medium (WNM) and dense cold neutral medium (CNM), believed to be maintained by two thermodynamical states of neutral hydrogen (Wolfire et al., 1995, 2003). Based on earlier studies, it has been argued that large amount of thermally unstable matter is not allowed to exist in theoretical models of global ISM (Kalberla & Haud, 2018) and even if exists, it has to be a transition phase. However, it has been subsequently found, in numerical simulations (Kritsuk et al., 2017; Kim et al., 2023) as well as observationally (Kalberla & Haud, 2018; Murray et al., 2018; Heiles & Troland, 2003; Roy et al., 2013) that a large fraction of the ISM is actually in the unstable neutral media (UNM), comprising about $\sim 20-40\%$, both in volume and mass fractions, making the ISM actually a three-phase medium. As such, we do not have sufficient physical understanding of the UNM phase, the presence which definitely affects the status of both WNM and CNM phases. We should note that besides these three phases, there is also a hot ionized phase, known as the hot ionized medium (HIM), existing at a much higher temperature, believed to be maintained by supernovae heating and a warm ionized phase, maintained by hot B-type stars, known as warm ionized medium (WIM) (McKee & Ostriker, 1977). We in this work, however, do not consider the WIM or HIM phase in our analysis.
@@
Starting from Field’s work (1965), there have been various one-dimensional (1D) studies of TI in both linear and nonlinear regimes with different background conditions. Hennebelle and others (1999; 2000) have observed that transonic converging flows of warm gas can produce long-lived (more than cooling time) condensations of cold gas and in the presence of a magnetic field, in which condensation forms along the field direction due to realignment of fluid flow if the angle between original flow and field is small enough. Another concurrent study of shock propagation into WNM and CNM observed fragmentation of thin dense, thermally collapsed layer into molecular clouds (Koyama & Inutsuka, 2000). They proceeded with several further simulations of TI in one, two, and three dimensions (Koyama & Inutsuka, 2000, 2002; Inutsuka, 2005). There have been many other studies which investigated propagation of a shock front in a thermally unstable gas driving two-phase turbulence in the ISM (Koyama & Inutsuka, 2002; Audit & Hennebelle, 2005), effect of kinematic viscosity on nonlinear turbulent fluid flow (Inutsuka, 2005; Koyama & Inutsuka, 2006), and nonlinear evolution of TI for converging flows of WNM in presence of a magnetic field (Inoue & Inutsuka, 2008, 2009). Koyama & Inutsuka (2004) included isotropic thermal conduction term and put forward the so-called ‘Field condition’ which limits the simulation cell size to avoid artificial fragmentation of the media. In a magnetized background, both pressure and thermal conduction become anisotropic, which leads to the formation of long filamentary condensations along the field direction (Choi & Stone, 2012). Recently, Jennings & Li (2021) have investigated the TI for multiphase gas in the ISM with a Gaussian random field (GRF) density perturbation, in the presence of realistic conduction. Apart from the ISM, there are several other studies on TI in ICM (Parrish et al., 2009; Sharma et al., 2010; McCourt et al., 2011) and solar corona regions (Soler et al., 2011; Antolin et al., 2022; Claes & Keppens, 2019). Falle et al. (2020) have reinvestigated Field’s linear theory of TI using modern wave interaction and stability analysis, providing a clearer representation of the underlying physical mechanisms. That being so, the study of TI with different initial conditions and new physical processes remains always an interesting area for understanding the balance of cooling and heating in the formation of multiphase structures in various astrophysical environments.
@@
We, in this work, perform an inviscid two-dimensional (2D) simulation of the TI in the ISM, with a background magnetic field with pressure anisotropy. We carry out our simulations through a multi-fluid, 2D flux-corrected transport code (mFCT), which takes care of any possible numerical diffusion. We use a polybaric pressure model in our simulation, where we handle the pressure anisotropy with an anisotropy parameter $a_{p}$ (Stasiewicz, 2004, 2005, 2007). With the polybaric pressure equation, we can model the ISM pressure as polytropic in nature, which is supported by observations (Spaans & Silk, 2000) in contrast to the widely used Chew-Goldberger-Low (CGL) double adiabatic pressure model (Chew et al., 1956), which requires very slow plasma flow so that the parallel and perpendicular pressures become completely disconnected, a condition very difficult to be justified (Kulsrud, 1983). The pressure anisotropy parameter is assumed to be constant which is also supported by the fact that variable pressure anisotropy would require the magnetic field to be strongly dependent on density, which however is not supported by observational data for the ISM density range $\sim 0.1-100\,{\rm cm}^{-3}$ (Troland & Heiles, 1986). The polybaric model has been successfully applied to study large amplitude magnetosonic oscillations in Earth’s magnetosheath region. As far as the kinematic viscosity is concerned, we can find many contemporary simulations with inviscid fluid (Padoan et al., 2016; Piontek & Ostriker, 2005; Sharma et al., 2010; Audit & Hennebelle, 2005; Soler et al., 2011; Inoue et al., 2006; Hennebelle & Audit, 2007; Kobayashi et al., 2022, 2020; Inoue & Inutsuka, 2012). We would also like to note that our analysis does not include any star formation activity (Kim et al., 2014, 2023) or supernova feedback term (Kim & Ostriker, 2017) which essentially requires external forcing. Our analysis is to assess the thermal evolution of the ISM and to see whether we can arrive at the similar outcome regarding distributions of WNM, UNM, and CNM with realistic pressure profile and time-dependent heating/cooling, without considering any external forcing terms. We further note that while it is now well-known that supernova feedback and other external forcing terms can drive shock fronts through the ISM; from simulation (Kim et al., 2014) and observational studies (Heiles & Troland, 2003) on $21\,\rm cm$ neutral hydrogen emission, it is observed that the so-called spin temperature and kinetic temperature remain almost equal in the CNM and UNM regions indicating near sonic or subsonic turbulent velocity, ruling out formation of any shock front, which is consistent with our analysis. As we shall see that the nature of the turbulence driven by thermal instability alone is quite similar in behaviour to those driven by external forcing (Kobayashi et al., 2022; Kritsuk et al., 2017; Hennebelle & Audit, 2007; Saury et al., 2014; Kim & Ostriker, 2017) such as converging shocks or supernova feedback, which indicates toward a universal nature of the turbulence, irrespective of being driven by external shock fronts or time-dependent heating and cooling.
@@ 2 Ideal MHD equations with radiative heat-loss
is a column vector whose components are mass density $\rho$, momentum density $\rho\bm{v}$, total energy density ${\cal E}$, and magnetic field $\bm{B}$. The quantity $\mathbb{F}$ represents the corresponding flux. The individual equations making up the above model are the continuity, momentum, energy equations, and Faraday’s law combined with Ohm’s law. The compressible fluid equations are given by
@@
$\rho\frac{d\bm{v}}{dt}$ $=$ $\bm{J}\times\bm{B}-\nabla\cdot\mathbb{P},$ (4)
$\frac{\partial{\cal E}}{\partial t}+\nabla\cdot\mathbb{F}_{\textrm{en}}$ $=$ $-\rho\mathfrak{L},$ (5)
@@
where $d/dt\equiv\partial/\partial t+\bm{v}\cdot\nabla$, $\mathbb{P}$ is the anisotropic pressure tensor, and
@@
The lone term on the right hand side of Eq.(5) is the radiative heat-loss term with $\mathfrak{L}$ as the total heat-loss function (cooling and heating) (Field, 1965). The total energy per unit volume ${\cal E}$ is given by
@@ 2.1 Generalized continuity equation form
The flux-conservative form of the above MHD equations can be expressed in a form of generalized continuity equation with a source term, which will actually be used in our computation. This form can be written as
@@
where we split up the total flux term into the flux of the quantity ${\cal F}$ and a corresponding source $S$. The column vector ${\cal F}$
${\cal F}=(\rho,\rho\bm{v},{\cal E},\bm{B})^{\prime}$ (15)
as before and the source terms are given by
@@
Note the above equations do not have the kinematic viscosity term.
@@ 2.2 Radiative heat-loss with thermal conduction
where the directions $\parallel$,$\perp$ , ⟂ refer to that of the equilibrium magnetic field $\bm{B}_{0}$. For interstellar matter (ISM), we follow a model for $K$ and $\mathfrak{L}$, which are prescribed as (Falle et al., 2020)
@@ 2.3 Equation of state
$\frac{d}{dt}\left(\frac{p_{\parallel}B^{2}}{\rho^{3}}\right)$ $=$ $0$ (30)
@@ 3.2 Normalization
We now normalize all the physical quantities by their respective equilibrium values (denoted with subscript ‘00’). Specifically, we adopt the following normalization
@@
For $T\sim T_{0}$, a quick calculation shows that $f_{1}\ll f_{2}$, appearing in the cooling functions Eqs.(25,26). So, in the normalized form, we can scale the radiative heat-loss term as
@@ 3.3 Numerical cycle
In the next sections, we shall present the results for the linear growth phase of the TI in a 1D case. The nonlinear simulations are run in 2D as mentioned in the previous section. In the 1D case, we use a sinusoidal perturbation, which helps us compare the linear growth phase of the TI with the analytical results (Field, 1965). In the 2D case, we use a small randomized density perturbation state with a perturbation level $\sim 1\%$ of the equilibrium density. The initial state is so chosen that it is slightly unstable to the thermal condensation mode. The use of energy equation is desirable for the purpose of conservation of the total energy. Once we determine the total energy, the $p_{\parallel,\perp}$ and the total pressure $p$ can be calculated using Eqs.(10,32). The temperature then can be determined from the scalar (averaged) pressure by using the normalised relation
@@
We use the FCT method to solve Eqs.(14-19) in a uniform grid of size $(1000\times 1000)$ with periodic boundary conditions.
@@ 3.4 The characteristic scales
where $K=K_{c}T^{1/2}$. The Field length defines the minimum scale that is required for the TI to grow or else, any density perturbation smaller than the Filed length will be quickly erased by the thermal conductivity. Though Koyama and Inutsuka (2000) have used a simplified version of the heat-loss function (Wolfire et al., 1995) to calculate $\lambda_{F}$, one can also use the heat-loss function as given by Eq.(23). For the parameters used in this work, the Field length comes out to be $\lambda_{F}\sim 0.24-0.35\,{\rm pc}$ for $T\sim 500-1500\,{\rm K}$. As we can see, our computational grid of cell size
$\Delta x=2\pi\times 10/1000\simeq 0.063\,{\rm pc}$
can very effectively resolve the Filed length. We also note that our time step size $\Delta t\lesssim 0.01$ is quite within the CFL limit

### v2

# Thermal instability and multiphase dynamics in the ISM with polybaric pressure effects

@@ Abstract
In this work, we have carried out a two-dimensional (2D) simulation of thermal instability (TI) in interstellar matter (ISM), considering it to be a weakly ionized inviscid plasma with radiation loss. We carry out the simulation using our multi-fluid flux-corrected transport (mFCT) code which incorporates a background magnetic field and anisotropic pressure. The anisotropic pressure is modeled with a polybaric pressure model. The findings of our analysis are consistent with the contemporary status of knowledge about the multiphase nature of the ISM with volume and mass fractions of the various components of the ISM that is warm, cold, and unstable neutral matter (UNM) in the ranges reported by various numerical and observational analysis. Though the strength of the background magnetic field only marginally affects the overall evolution, the ratio of the parallel and perpendicular pressures can considerably alter the mass and volume fractions of the three phases, which can affect the overall evolution of the TI in the long run.
keywords:
instabilities – (magnetohydrodynamics) MHD – plasmas – ISM: clouds – turbulence – methods: numerical
pubyear: 2024pagerange: Thermal instability and multiphase dynamics in the ISM with polybaric pressure effectsThermal instability and multiphase dynamics in the ISM with polybaric pressure effects
@@ 1 Introduction
Thermal instability (TI) is a physical process responsible for non-gravitational condensations in various plasma environments such as solar corona, solar wind, magnetosheath, interstellar medium (ISM), intracluster medium (ICM) etc. When thermal equilibrium becomes unstable, it can change the structure and dynamics of the medium at a small scale resulting in condensations of high density and low temperature such as clouds and filaments, known as the precursors to star formation. First introduced by Parker (1953) and then by Field in his seminal work (1965), TI is known to explain the existence of the multiphase structure of the ISM and mass flow between different intermediate structures, which has very significant effects on the rate of star formation (Chieze, 1987).
@@
The multiphase model of the ISM is a well-studied physical model, which was thought to be consisting of two stable phases – warm and cold, formed out of neutral atomic hydrogen (Field et al., 1969). Studies have later confirmed the existence of these two phases – diffuse warm neutral medium (WNM) and dense cold neutral medium (CNM), believed to be maintained by two thermodynamical states of neutral hydrogen (Wolfire et al., 1995, 2003). Based on earlier studies, it has been argued that large amount of thermally unstable matter is not allowed to exist in theoretical models of global ISM (Kalberla & Haud, 2018) and even if exists, it has to be a transition phase. However, it has been subsequently found, in numerical simulations (Kritsuk et al., 2017; Kim et al., 2023) as well as observationally (Kalberla & Haud, 2018; Murray et al., 2018; Heiles & Troland, 2003; Roy et al., 2013) that a large fraction of the ISM is actually in the unstable neutral media (UNM), comprising about $\sim 20-40\%$, both in volume and mass fractions, making the ISM actually a three-phase medium. As such, we do not have sufficient physical understanding of the UNM phase, the presence of which definitely affects the status of both WNM and CNM phases. We should note that besides these three phases, there is also a hot ionized phase, known as the hot ionized medium (HIM), existing at a much higher temperature, believed to be maintained by supernovæ heating and a warm ionized phase, maintained by hot B-type stars, known as warm ionized medium (WIM) (McKee & Ostriker, 1977). We in this work however, do not consider the WIM or HIM phase in our analysis.
@@
Starting from Field’s work (1965), there have been various one-dimensional (1D) studies of TI in both linear and nonlinear regimes with different background conditions. Hennebelle and others (1999; 2000) have observed that transonic converging flows of warm gas can produce long-lived (more than cooling time) condensations of cold gas and in the presence of a magnetic field, in which condensation forms along the field direction due to realignment of fluid flow if the angle between original flow and field is small enough. Another concurrent study of shock propagation into WNM and CNM observed fragmentation of thin dense, thermally collapsed layer into molecular clouds (Koyama & Inutsuka, 2000). They proceeded with several further simulations of TI in one, two, and three dimensions (Koyama & Inutsuka, 2000, 2002; Inutsuka, 2005). There have been many other studies which investigated propagation of a shock front in a thermally unstable gas driving two-phase turbulence in the ISM (Koyama & Inutsuka, 2002; Audit & Hennebelle, 2005), effect of kinematic viscosity on nonlinear turbulent fluid flow (Inutsuka, 2005; Koyama & Inutsuka, 2006), and nonlinear evolution of TI for converging flows of WNM in presence of a magnetic field (Inoue & Inutsuka, 2008, 2009). Koyama & Inutsuka (2004) included isotropic thermal conduction term and put forward the so-called ‘Field condition’ which limits the simulation cell size to avoid artificial fragmentation of the medium. In a magnetized background, both pressure and thermal conduction become anisotropic, which leads to the formation of long filamentary condensations along the field direction (Choi & Stone, 2012). Recently, Jennings & Li (2021) have investigated the TI for multiphase gas in the ISM with a Gaussian random field (GRF) density perturbation, in the presence of realistic conduction. Apart from the ISM, there are several other studies on TI in ICM (Parrish et al., 2009; Sharma et al., 2010; McCourt et al., 2011) and solar corona regions (Soler et al., 2011; Antolin et al., 2022; Claes & Keppens, 2019). Falle et al. (2020) have reinvestigated Field’s linear theory of TI using modern wave interaction and stability analysis, providing a clearer representation of the underlying physical mechanisms. That being so, the study of TI with different initial conditions and new physical processes remain always an interesting area for understanding the balance of cooling and heating in the formation of multiphase structures in various astrophysical environments.
@@
We, in this work perform an inviscid two-dimensional (2D) simulation of the TI in the ISM, with a background magnetic field with pressure anisotropy. We carry out our simulations through a multi-fluid, 2D flux-corrected transport code (mFCT), which takes care of any possible numerical diffusion. We use a polybaric pressure model in our simulation, where we handle the pressure anisotropy with an anisotropy parameter $a_{p}$ (Stasiewicz, 2004, 2005, 2007). With the polybaric pressure equation, we can model the ISM pressure as polytropic in nature, which is supported by observations (Spaans & Silk, 2000) in contrast to the widely used Chew-Goldberger-Low (CGL) double adiabatic pressure model (Chew et al., 1956), which requires very slow plasma flow so that the parallel and perpendicular pressures become completely disconnected, a condition very difficult to be justified (Kulsrud, 1983). The pressure anisotropy parameter is assumed to be constant which is also supported by the fact that variable pressure anisotropy would require the magnetic field to be strongly dependent on density, which however is not supported by observational data for the ISM density range $\sim 0.1-100\,{\rm cm}^{-3}$ (Troland & Heiles, 1986). The polybaric model has been successfully applied to study large amplitude magnetosonic oscillations in Earth’s magnetosheath region. As far as the kinematic viscosity is concerned, we can find many contemporary simulations with inviscid fluid (Padoan et al., 2016; Piontek & Ostriker, 2005; Sharma et al., 2010; Audit & Hennebelle, 2005; Soler et al., 2011; Inoue et al., 2006; Hennebelle & Audit, 2007; Kobayashi et al., 2022, 2020; Inoue & Inutsuka, 2012). We also note that supernovæ activity often provides the essential triggering mechanism which can drive the thermal instability. We incorporate the supernova feedback (or supernova forcing) activity (Kim & Ostriker, 2017) through a random forcing term in our equations. However, we must emphasize that we are only interested in large-scale structure of the ISM and assume that supernovæ shock fronts have decayed into wide-scale random perturbation with solenoidal and compressional components (Padoan et al., 2016). Our analysis is to primarily assess the effect of polybaric pressure on thermal evolution of the ISM and to see whether we can arrive at the similar outcome regarding distributions of WNM, UNM, and CNM with time-dependent heating/cooling. We further note that while supernova feedback can drive shock fronts through the ISM; from simulation (Kim et al., 2014) and observational studies (Heiles & Troland, 2003) on $21\,\rm cm$ neutral hydrogen emission, it is observed that the so-called spin temperature and kinetic temperature remain almost equal in the CNM and UNM regions indicating near sonic or subsonic turbulent velocity, ruling out existence of any shock front, which is consistent with our assumption (Kobayashi et al., 2022; Kritsuk et al., 2017; Hennebelle & Audit, 2007; Saury et al., 2014; Kim & Ostriker, 2017).
@@ 2 Ideal MHD equations with radiative heat-loss
[…]
@@
$\frac{\partial}{\partial t}\left(\rho\bm{v}\right)+\nabla\cdot\left(\rho\bm{v}\bm{v}\right)$ $=$ $\bm{J}\times\bm{B}-\nabla\cdot\mathbb{P}+{\bm{f}}_{\rm SN},$ (4)
$\frac{\partial{\cal E}}{\partial t}+\nabla\cdot\mathbb{F}_{\textrm{en}}$ $=$ ${\bm{v}}\cdot{\bm{f}}_{\rm SN}-\rho\mathfrak{L},$ (5)
@@
where $\mathbb{P}$ is the anisotropic pressure tensor, and
@@
[…]
@@
where $\mu$ is the mean molecular weight. In the above expression, $\hat{b}_{i}$ are the components of the magnetic field vector $(\bm{B}/B)$. For the combined Faraday’s and Ohm’s law, we have
$\mathbb{F}_{B}=\bm{v}\bm{B}-\bm{B}\bm{v}.$ (13)
@@ 2.1 Generalized continuity equation form
To obtain the numerical results presented in this paper, we have employed the Flux-Corrected-Transport (FCT) scheme (Section 3.1), which is based upon conservation of fluxes while transporting any fluid quantity across structured grids by suppressing numerical diffusion that may arise in the system from discretization of the differential operators. Implementation of the FCT formalism requires the fluid equations to be put in the form of generalized continuity equations. The flux-conservative form of the above MHD equations can be expressed in a form of generalized continuity equation with a source term, which will actually be used in our computation. This form can be written as
@@
[…]
@@
Note that the supernova forcing terms are not included in the above expressions for brevity and and can always be added afterwards.
@@ 2.2 Radiative heat-loss with thermal conduction
where the directions $\parallel,\perp$ refer to that of the equilibrium magnetic field $\bm{B}_{0}$. For interstellar matter (ISM), we follow a model for $K$ and $\mathfrak{L}$, which are prescribed as (Falle et al., 2020)
@@ 2.3 Equation of state
$\frac{d}{dt}\left(\frac{p_{\parallel}B^{2}}{\rho^{3}}\right)$ $=$ $0.$ (29)
@@ 3.2 Normalization
We now normalize all the physical quantities by their respective equilibrium values (denoted with subscript ‘0’). Specifically, we adopt the following normalization
@@
For $T\sim T_{0}$, a quick calculation shows that $f_{1}\ll f_{2}$, appearing in the cooling functions Eqs.(24,25). So, in the normalized form, we can scale the radiative heat-loss term as
@@ 3.3 Numerical cycle
The nonlinear simulations are run in 2D as mentioned in the previous section, where we use the supernova forcing term ${\bm{f}}_{\rm SN}$ to drive the perturbation through a randomised velocity field with a mix of solenoidal and compressive components. Note that the initial state is unstable to the thermal condensation mode. The use of energy equation is desirable for the purpose of conservation of the total energy. Once we determine the total energy, the $p_{\parallel,\perp}$ and the total pressure $p$ can be calculated using Eqs.(10,31). The temperature then can be determined from the scalar (averaged) pressure by using the normalised relation
@@
We use the FCT method to solve Eqs.(14-18) in a uniform grid of size $(1000\times 1000)$ with periodic boundary conditions.
In Section 4, we present the results for the linear growth phase of the TI in a 1D case. In the 1D case, we use a sinusoidal perturbation, which helps us compare the linear growth phase of the TI with the analytical results (Field, 1965).
@@ 3.4 The characteristic scales
where $K=K_{c}T^{1/2}$. The Field length defines the minimum scale that is required for the TI to grow or else, any density perturbation smaller than the Field length will be quickly erased by the thermal conductivity. Though Koyama and Inutsuka (2000) used a simplified version of the heat-loss function (Wolfire et al., 1995) to calculate $\lambda_{F}$, one can also use the heat-loss function as given by Eq.(22). For the parameters used in this work, the Field length comes out to be $\lambda_{F}\sim 0.24-0.35\,{\rm pc}$ for $T\sim 500-1500\,{\rm K}$. As we can see, our computational grid of cell size
$\Delta x=2\pi\times 10/1000\simeq 0.063\,{\rm pc}$ (50)
can very effectively resolve the Field length. We also note that our time step size $\Delta t\lesssim 0.01$ is quite within the CFL limit
@@
At this point, we note that the Field length, as determined from the linear dispersion relation can be written in terms of the characteristic wavenumber of the isobaric perturbation due to radiation condensations divided by the average mean free path of conducting particles (Field, 1965)
$k_{F}^{2}=k_{K}(k_{T}-k_{\rho}),$ (57)
where $k_{F}$ is the wavenumber corresponding to the Field length, $k_{T,\rho}$ are the wavenumbers of isothermal and isochoric perturbations, and $k_{K}$ is the reciprocal of the mean free path of the conducting particles. The corresponding Field length is then given by
$\lambda_{F}\sim k_{F}^{-1}=\left[\frac{\rho_{0}}{K}\left(\mathfrak{L}_{T}-\frac{\rho_{0}}{T_{0}}\mathfrak{L}_{\rho}\right)\right]^{-1/2}.$ (58)
The relation given by Eq.(49) (Koyama & Inutsuka, 2004) is actually obtained by considering isochoric perturbation
$\lambda_{F}\sim\left(\frac{K}{\rho_{0}\mathfrak{L}_{T}}\right)^{1/2}$ (59)
and approximating $\mathfrak{L}=\left(\rho_{0}/m_{H}^{2}\right)\Lambda(T)$ and writing $\mathfrak{L}_{T}\sim\mathfrak{L}/T$. As we can see that the Field length obtained from linear dispersion relation Eq.(58) diverges near the marginal stability point where $\mathfrak{L}_{T,\rho}\to 0$. As such, it should not be interpreted as a literal physical scale in the nonlinear regime, but rather as a guideline for the smallest linearly unstable mode and as a numerical resolution criterion (Koyama & Inutsuka, 2004). In our nonlinear simulations, characteristic condensation scales are instead set by the interplay of turbulence and pressure balance. We can apply the same argument for the cooling as well as conduction lengths.
3.5 Grid convergence
The FCT method with Zalesak’s flux limiter (Zalesak, 1979) enforces monotonicity by restricting anti-diffusive fluxes. But, it can still retain some numerical diffusion in the process which may cause the solution to evolve into a staircase of square-wave like structure. However, with sufficient grid resolution, this effect can be minimized keeping the underlying physics intact.
Figure 1: (a) The temporal density evolution at different grid size is shown. Note that a grid size of $1000$ corresponds to a grid resolution of $0.063\,{\rm pc}$, which is about 10 times smaller than the so-called Field length. The legends in the figure indicate the grid sizes and they all overlap within the accuracy limit. (b) Relative density evolution $(\delta\rho/\delta\rho_{1000})$ in time, where $\delta\rho_{1000}$ is the change in density with a grid size of $1000$. The shaded region in the figure shows $\pm 1\%$ band around the $\delta\rho_{1000}$ value. We see that relative density for all the grid size (from $250$ onwards till $1000$) fall within this $\pm 1\%$ band, indicating a grid-independent evolution.
To check the grid-dependence in our simulation, in Fig.1, we have shown the temporal evolution of the absolute change in normalised density (left), which corresponds to Fig.2(d) later in Section 4, and relative density (right) are shown for linear stage of the instability. In Fig.1(a), the temporal density evolution for grid sizes $250,500,600,750,800$, and $1000$ are shown, which all overlap with each other. The highest grid resolution is for grid size $1000$, which corresponds to a grid resolution of $0.063\,{\rm pc}$ that is about $10$ times smaller than the so-called Field length and the lowest grid size $250$ corresponds to resolution of $0.252\,{\rm pc}$, which is just about the Field length. The legends in the figure show various grid sizes. The right panel shows the relative density deviations ($\delta\rho/\delta\rho_{1000}$) for different grid sizes ($\delta\rho_{1000}$ being the density for grid size $1000$) and the shaded region indicates $\pm 1\%$ band around the $\delta\rho_{1000}$. As we can see that all the density values are within the $\pm 1\%$, indicating grid-convergence of the simulation results.